# Revolutionizing Future Connectivity: A Contemporary Survey on AI-empowered Satellite-based Non-Terrestrial Networks in 6G

### v3

@@ III-D1 Supervised Learning (SL)
In an SL model, a training dataset containing a set of features as inputs and corresponding current outputs is provided to the model. The model with an initial set of parameters is trained through a number of iterations for mapping inputs to outputs. As the output label is clearly defined, the model can improve its performance by comparing its predicted outputs with the actual outputs [40]. SL problems can be broadly classified into two categories depending on the type of output labels: regression and classification problems. Regression [41] is a statistical method that investigates the relationship between a dependent (target) variable to one or more independent (given) variables. In this method, the functional mapping between inputs and outputs is estimated by minimizing the error between the predicted and actual outputs. Here the output label can be continuous. In classification, the output labels correspond to distinct classes arising in computer vision, image classification, etc. Generally, classification problems are solved by using probabilistic classifiers to map output classes from inputs. To train complex SL problems, NNs are used to learn complicated functional mapping between inputs and outputs. We discuss the major ML and DL approaches in the context of SL problems below:
@@
ML Approaches: There are a number of SL algorithms to train the model. Linear regression [42] focuses on regression problems, whereas logistic regression [43] focuses on classification problems. Decision tree is used in classification problems by forming a tree-like structure to learn the best split at every node level based on a statistical measure like information gain [44]. The classification starts at the root node and traverses down along the branches based on intermediate decisions till the leaf nodes which represent the final classification decisions. Naive Bayes Model [45] is a form of a simple probabilistic classifier that uses the Bayesian Theorem to decide the classes under the strong assumption of feature independence. It is very useful, especially in high-dimensional classification problems. Support Vector Machine (SVM) [46] is another important type of classifier that decides the splitting hyperplane between different classes by maximizing the distances between the nearest data point (in both classes) and the hyperplane.
@@
DL Approaches: Different DL approaches are also proposed in the literature to tackle complicated SL problems effectively. Perceptron [47] is one of the first NN architectures that have been proposed. It is a single-layer NN that can do binary classification like logistic regression. The main difference is to introduction of a simple activation function (step function) as a first step to more complex and advanced architectures. The simplest multi-layer NN architecture is the Fully Connected Neural Networks (FCNN) (Figure 8). This is also known as Multi-Layer Perceptron (MLP). It has multiple hidden layers between the input and output layers without any back loops. As the name suggests, all the neurons between two adjacent layers are connected to each other. Extreme Learning Machine (ELM) [48] is a very special type of NNs where the neurons are randomly connected and the training is done one-shot using least square fits. Another different type of NN is the Deep Residual Network (DRN) [49] with extra connections passing input from one layer to a later layer as well as the next layer. There are also Probabilistic Neural Networks (PNN) [50] which can recognize the underlying pattern and generate the probability distribution function for different classes.
@@
Convolutional Neural Networks (CNN) [51] is an important type of NN that can take multidimensional inputs like images and classify them with great accuracy by discovering spatial features (Figure 8). The CNNs are composed of convolutional layers and subsequent pooling layers. The convolutional layers divide the whole input into smaller blocks and scan through them to learn the different features. The idea is to exploit the high correlation among neighboring cells with reduced complexity. A pooling layer is used to simplify this extraction process by getting rid of redundant features. Often CNN is accompanied by an FCNN to take care of nonlinearity and generate final classification results. A counterpart of CNN is the Deconvolutional Network (DN) [52] which takes the classes as inputs and generates CNN input features by comparing them with actual CNN inputs.
@@
Another important type of NN is Recurrent Neural Networks (RNN) [53] with back loops. So the neurons in a layer are not only connected to previous layer neurons but also can be connected to the neurons from the subsequent layers. (Figure 10) This allows it to capture temporal correlation among different layers and can be useful where decisions from past iterations or samples can influence current ones. However, they suffer from vanishing gradient issues due to long-term temporal dependencies [54]. To tackle this issue more sophisticated architectures like Gated recurrent units (GRU) [55] and Long-Term Short Memory (LSTM) [56] with special memory cells and gates are introduced. Reservoir computing (RC) [57] is a low training complexity RNN framework for computation where the inputs are fed into a fixed and non-linear system, known as a reservoir, and then mapped into outputs from the reservoir neurons. Liquid State Machines (LSMs) are examples of RCs where the neurons are randomly connected receiving time-varying inputs. Echo-state networks (ESNs) are also a type of RC that uses a sparsely connected hidden layer (reservoir) with typically 1% connectivity. The connectivity and weights of hidden neurons are fixed and randomly assigned.
@@
Another significant advancement in deep learning architecture, known as transformers, holds immense promise in the development of intelligent systems, particularly in communication environments. The transformer is a sequence-to-sequence neural network model comprising both an encoder and a decoder module, each with an identical architecture [58]. To streamline the input and output sequences, embedding and positional encoding layers are employed. Both the encoder and decoder primarily consist of a self-attention sub-layer and a position-wise sub-layer, with an additional masked attention sub-layer in the decoder. Each sub-layer is complemented by a residual connection and normalization module, facilitating the capture of long-range dependencies within the input data through self-attention.
@@ III-D2 Unsupervised Learning (UL)
In UL, a raw unlabeled dataset is provided to discover existing patterns and features [59] using some statistical learning approach. This is very useful when the data is not labeled. The algorithms find the underlying structure of the data and predict the outputs by adapting the model. Here the classes are not explicitly stated, so the classes need to be generated based on the distribution of input features in multi-dimensional spaces. It can be even used for generating labeled data to transform the original problem into an SL problem, which is usually easier to solve. Furthermore, clustering is another important UL problem where the model outputs different clusters based on the inherent pattern of data distribution. Dimensionality reduction can be also classified as a UL problem as it reduces the state space of the feature vectors in a general ML setup.
@@
ML Approaches: There are a number of unsupervised learning algorithms in the literature. Principal Component Analysis (PCA) [60] is primarily used for dimensionality reduction of a high dimensional dataset. It reduces the number of correlated features converting them into a set of uncorrelated features, which are also termed principal components, using orthogonal transformation of basis vectors. Reducing the dimensions of inputs also reduces the number of features to be learned, which later can be leveraged in SL techniques. It is sometimes not considered an UL technique, but rather a preprocessing technique for data analysis with reduced dimensions. In Probabilistic Graph Models (PGMs), the probabilistic relationship between random variables is modeled through a graph [61].
@@
K-means Clustering [62] divides all the data points into K clusters in which each data point belongs to the cluster having the nearest mean. The mean of the data points in a particular cluster defines the center of the cluster. Another variant of K-means Clustering is called K-medoids Clustering where the centralmost data point of a cluster is defined as the center of the cluster [63]. Various mixture models, such as the finite mixture model, Gaussian Mixture Model (GMM) [64], etc. are also used for clustering. Hierarchical clustering can cluster data into a hierarchy of groups without predefining the number of clusters. It also comes with increasing computational costs compared to other clustering approaches. k-Nearest Neighbours (KNN) [65] algorithm determines the k-nearest neighbors for all the data points of an unknown feature vector whose class is to be identified.
@@
DL Approaches: Generally speaking, autoencoders [66] are used to help reduce the noise in data. In an autoencoder, first, we encode a high dimensional input, then decode it to reconstruct the input at the output again (Figure 10). The intermediate hidden layer neurons represent a compressed representation of the inputs getting rid of irrelevant and noisy components. Some other variations of this architecture are variational [67], noisy [68], and sparse [69] autoencoders. In variational autoencoders, the compact representation of data is used to generate new data points sampling from the latent space. In sparse autoencoders, the loss function is also expanded by adding another term called sparsity penalty regularization term for encouraging sparsity in the learned representations. In denoising autoencoders, robust representations are learned from the noisy input data.
@@
Deep Belief Networks (DBNs) [70] is a probabilistic generative graph model composed of hierarchical layers representing feature vectors. Here the top layers create undirected symmetric connections among them forming an associative memory. Greedy layer-wise training can be used for DBNs [71]. Two symmetric DBNs can be extended to the structure of deep autoencoders for efficiently decoding the feature vectors [72]. To use the feature extraction capability CNNs for UL, a combination of CNN and DBN is used in [73]. Hopfield NN [74] is a cyclic recurrent NN architecture where all the nodes are connected to each other. This provides an abstraction of circular shift register memory to form a global energy function and finding clusters without a supervisor. The Boltzmann Machine is another type of recurrent NN that has a stochastic symmetric recurrent architecture [75]. As the convergence rate is generally slow for these NNs, a variant of this, Restricted Boltzmann Machine (RBM) is designed to learn the probability distribution over input data but in a layered manner [76].
@@
In UL, competitive learning approaches, such as Self-Organizing Maps (SOMs) [77], each neuron competes to represent an input subset. Here a single neuron from a group of output neurons is activated while the other neurons adjust their individual values in regard to input data distribution. Generative Adversarial Networks (GAN) [78] consist of any two networks, with one generating data (generative network) and the other judging the generated data (discriminating network). The prediction accuracy of the discriminating network is then used to evaluate the error for the generating network. This creates a form of competition between the discriminator and the generator to get better in their corresponding tasks. We can also use ensemble learning methods [79] comprising multiple learning methods for better performances.
@@
Generative Diffusion Models (GDMs), as introduced in [80], represents a recent breakthrough in UL leveraging DL techniques, drawing inspiration from the principles of thermodynamic diffusion. GDMs have gained widespread recognition due to their remarkable ability to generate high-quality data and simple implementation procedures. In contrast to GANs, GDMs employ a denoising network that iteratively converges to an estimate of the real sample. This model works in two distinct phases: the forward and reverse diffusion processes [81]. In the forward diffusion phase, Gaussian noise is gradually introduced through a series of steps to create the target input for the denoising network. Subsequently, the denoising network is trained to reverse the noise effect for generating the original content.
@@ III-D3 Reinforcement Learning (RL)
In RL, an agent learns to behave in a particular environment by performing thousands of actions and getting rewards or penalties based on those actions [82]. This behavior (formally known as policy) is defined by the set of actions the agent learns from its experiences. The environment is defined by some mathematical models, the most common one is the Markov Decision Process (MDP) [83]. Here the feedback is neither provided using explicit labels like in SL nor the model is learned like in UL, but the behavior of an agent is learned through the rewards or penalties based on the set of actions taken by going from one state to another with a transition probability. The goal is to find out the optimum policy so that the total reward can be maximized (or the total penalty can be minimized) over a horizon of future time intervals given the current state of the agent. In Figure 11, we show a generic structure of an RL framework as an MDP model.
@@
ML Approaches: Depending on whether an RL model is explicitly created or not, RL can be fundamentally divided into two major categories: model-based and model-free RL methods [84]. In the model-based methods, the transition probabilities between different states are assumed to be known, whereas, in model-free methods, these probabilities are learned through iterations. Dynamic Programming (DP) [85] is the most popular model-based method used in practice. However, model-free methods like the Monte Carlo (MC) method [86] are most commonly used due to their flexibility and practicality in real systems. Q-learning [87] is one of the most popular model-free methods where ”Q” refers to the expected rewards for an action taken in a given state over the time horizon, known as the value function. Another important counterpart of Q-learning is the State-Action-Reward-State-Action (SARSA) learning method where the agent learns the optimal policy in an online fashion [88]. This Q-learning is extended to the context of stochastic games [89] involving multiple agents in [90], which is also known as the Multi-Agent Reinforcement Learning (MARL) method. Another approach for model-free RL is to learn the policy directly instead of learning the value functions, which is known as the Policy Gradient (PG) method [91]. If the policy gradient is estimated in a deterministic fashion, it is called the Deterministic Policy Gradient (DPG) method [92]. To have the benefits of both approaches, the Actor-Critic (AC) method is proposed in [93] where the critic estimates the value function and the actor updates the policy gradient in the direction suggested by the critic.
@@
DL Approaches: General RL approaches without NNs can work well for small-scale experiments. However, when the action or state space is really large, the computation complexity exponentially increases. This phenomenon is quite common in practical systems like communication networks. To estimate the value functions or policies with RL frameworks with large state or action spaces, DL approaches can be very useful. This learning approach is also known as Deep Reinforcement Learning (DRL). The most popular and simplest DRL approach is the Deep Q-Network (DQN) [94] method. In this approach, instead of an iterative approach for updating Q values in the Q-table, an NN is used to estimate the Q function value approximately. To prevent a large overestimation of action values, another DL framework is introduced on top of DQN for a fair evaluation of policies in Double Deep-Q-Network (DDQN) [95, 96]. In Dueling Deep-Q-Network [97] method, both state and action values are separately estimated. As the expected value function may be overestimated as the expected value does not capture the complete probability distribution of random variables, Distributional Deep-Q-Network [98] is considered to update the Q function value based on its distribution. In the case of continuous action spaces, DL aided DPG method, known as Deep Deterministic Policy Gradient (DDPG) Q-Learning [99] provides better results. To deal with partial observable environments, Deep Recurrent Q-Networks (DRQN) [100] by introducing an LSTM layer in the FCNN architecture of DQN. Similarly, as for RL, MARL approaches can be efficiently solved using DRL architecture for each agent, namely known as a Multi-Agent Deep Reinforcement Learning (MADRL) framework.

### v4

@@ III-D1 Supervised Learning (SL)
In an SL model, a training dataset containing a set of features as inputs and corresponding current outputs is provided to the model. The model with an initial set of parameters is trained through a number of iterations for mapping inputs to outputs. As the output label is clearly defined, the model can improve its performance by comparing its predicted outputs with the actual outputs [ai-sl]. SL problems can be broadly classified into two categories depending on the type of output labels: regression and classification problems. Regression [reg] is a statistical method that investigates the relationship between a dependent (target) variable to one or more independent (given) variables. In this method, the functional mapping between inputs and outputs is estimated by minimizing the error between the predicted and actual outputs. Here the output label can be continuous. In classification, the output labels correspond to distinct classes arising in computer vision, image classification, etc. Generally, classification problems are solved by using probabilistic classifiers to map output classes from inputs. To train complex SL problems, NNs are used to learn complicated functional mapping between inputs and outputs. We discuss the major ML and DL approaches in the context of SL problems below:
@@
ML Approaches: There are a number of SL algorithms to train the model. Linear regression [linreg] focuses on regression problems, whereas logistic regression [logreg] focuses on classification problems. Decision tree is used in classification problems by forming a tree-like structure to learn the best split at every node level based on a statistical measure like information gain [dec-tree]. The classification starts at the root node and traverses down along the branches based on intermediate decisions till the leaf nodes which represent the final classification decisions. Naive Bayes Model [naive] is a form of a simple probabilistic classifier that uses the Bayesian Theorem to decide the classes under the strong assumption of feature independence. It is very useful, especially in high-dimensional classification problems. Support Vector Machine (SVM) [svm] is another important type of classifier that decides the splitting hyperplane between different classes by maximizing the distances between the nearest data point (in both classes) and the hyperplane.
@@
DL Approaches: Different DL approaches are also proposed in the literature to tackle complicated SL problems effectively. Perceptron [perceptron] is one of the first NN architectures that have been proposed. It is a single-layer NN that can do binary classification like logistic regression. The main difference is to introduction of a simple activation function (step function) as a first step to more complex and advanced architectures. The simplest multi-layer NN architecture is the Fully Connected Neural Networks (FCNN) (Figure 8). This is also known as Multi-Layer Perceptron (MLP). It has multiple hidden layers between the input and output layers without any back loops. As the name suggests, all the neurons between two adjacent layers are connected to each other. Extreme Learning Machine (ELM) [elm] is a very special type of NNs where the neurons are randomly connected and the training is done one-shot using least square fits. Another different type of NN is the Deep Residual Network (DRN) [drn] with extra connections passing input from one layer to a later layer as well as the next layer. There are also Probabilistic Neural Networks (PNN) [pnn] which can recognize the underlying pattern and generate the probability distribution function for different classes.
@@
Convolutional Neural Networks (CNN) [cnn] is an important type of NN that can take multidimensional inputs like images and classify them with great accuracy by discovering spatial features (Figure 8). The CNNs are composed of convolutional layers and subsequent pooling layers. The convolutional layers divide the whole input into smaller blocks and scan through them to learn the different features. The idea is to exploit the high correlation among neighboring cells with reduced complexity. A pooling layer is used to simplify this extraction process by getting rid of redundant features. Often CNN is accompanied by an FCNN to take care of nonlinearity and generate final classification results. A counterpart of CNN is the Deconvolutional Network (DN) [dn] which takes the classes as inputs and generates CNN input features by comparing them with actual CNN inputs.
@@
Another important type of NN is Recurrent Neural Networks (RNN) [rnn] with back loops. So the neurons in a layer are not only connected to previous layer neurons but also can be connected to the neurons from the subsequent layers. (Figure 10) This allows it to capture temporal correlation among different layers and can be useful where decisions from past iterations or samples can influence current ones. However, they suffer from vanishing gradient issues due to long-term temporal dependencies [vg]. To tackle this issue more sophisticated architectures like Gated recurrent units (GRU) [gru] and Long-Term Short Memory (LSTM) [lstm] with special memory cells and gates are introduced. Reservoir computing (RC) [rc] is a low training complexity RNN framework for computation where the inputs are fed into a fixed and non-linear system, known as a reservoir, and then mapped into outputs from the reservoir neurons. Liquid State Machines (LSMs) are examples of RCs where the neurons are randomly connected receiving time-varying inputs. Echo-state networks (ESNs) are also a type of RC that uses a sparsely connected hidden layer (reservoir) with typically 1% connectivity. The connectivity and weights of hidden neurons are fixed and randomly assigned.
@@
Another significant advancement in deep learning architecture, known as transformers, holds immense promise in the development of intelligent systems, particularly in communication environments. The transformer is a sequence-to-sequence neural network model comprising both an encoder and a decoder module, each with an identical architecture [trans]. To streamline the input and output sequences, embedding and positional encoding layers are employed. Both the encoder and decoder primarily consist of a self-attention sub-layer and a position-wise sub-layer, with an additional masked attention sub-layer in the decoder. Each sub-layer is complemented by a residual connection and normalization module, facilitating the capture of long-range dependencies within the input data through self-attention.
@@ III-D2 Unsupervised Learning (UL)
In UL, a raw unlabeled dataset is provided to discover existing patterns and features [ai-ul] using some statistical learning approach. This is very useful when the data is not labeled. The algorithms find the underlying structure of the data and predict the outputs by adapting the model. Here the classes are not explicitly stated, so the classes need to be generated based on the distribution of input features in multi-dimensional spaces. It can be even used for generating labeled data to transform the original problem into an SL problem, which is usually easier to solve. Furthermore, clustering is another important UL problem where the model outputs different clusters based on the inherent pattern of data distribution. Dimensionality reduction can be also classified as a UL problem as it reduces the state space of the feature vectors in a general ML setup.
@@
ML Approaches: There are a number of unsupervised learning algorithms in the literature. Principal Component Analysis (PCA) [pca] is primarily used for dimensionality reduction of a high dimensional dataset. It reduces the number of correlated features converting them into a set of uncorrelated features, which are also termed principal components, using orthogonal transformation of basis vectors. Reducing the dimensions of inputs also reduces the number of features to be learned, which later can be leveraged in SL techniques. It is sometimes not considered an UL technique, but rather a preprocessing technique for data analysis with reduced dimensions. In Probabilistic Graph Models (PGMs), the probabilistic relationship between random variables is modeled through a graph [pgm].
@@
K-means Clustering [kmean] divides all the data points into K clusters in which each data point belongs to the cluster having the nearest mean. The mean of the data points in a particular cluster defines the center of the cluster. Another variant of K-means Clustering is called K-medoids Clustering where the centralmost data point of a cluster is defined as the center of the cluster [kmed]. Various mixture models, such as the finite mixture model, Gaussian Mixture Model (GMM) [gmm], etc. are also used for clustering. Hierarchical clustering can cluster data into a hierarchy of groups without predefining the number of clusters. It also comes with increasing computational costs compared to other clustering approaches. k-Nearest Neighbours (KNN) [knn] algorithm determines the k-nearest neighbors for all the data points of an unknown feature vector whose class is to be identified.
@@
DL Approaches: Generally speaking, autoencoders [ae] are used to help reduce the noise in data. In an autoencoder, first, we encode a high dimensional input, then decode it to reconstruct the input at the output again (Figure 10). The intermediate hidden layer neurons represent a compressed representation of the inputs getting rid of irrelevant and noisy components. Some other variations of this architecture are variational [ae1], noisy [ae2], and sparse [ae3] autoencoders. In variational autoencoders, the compact representation of data is used to generate new data points sampling from the latent space. In sparse autoencoders, the loss function is also expanded by adding another term called sparsity penalty regularization term for encouraging sparsity in the learned representations. In denoising autoencoders, robust representations are learned from the noisy input data.
@@
Deep Belief Networks (DBNs) [dbn1] is a probabilistic generative graph model composed of hierarchical layers representing feature vectors. Here the top layers create undirected symmetric connections among them forming an associative memory. Greedy layer-wise training can be used for DBNs [dbn2]. Two symmetric DBNs can be extended to the structure of deep autoencoders for efficiently decoding the feature vectors [ae-dbn]. To use the feature extraction capability CNNs for UL, a combination of CNN and DBN is used in [cnn-dbn]. Hopfield NN [hnn] is a cyclic recurrent NN architecture where all the nodes are connected to each other. This provides an abstraction of circular shift register memory to form a global energy function and finding clusters without a supervisor. The Boltzmann Machine is another type of recurrent NN that has a stochastic symmetric recurrent architecture [bm]. As the convergence rate is generally slow for these NNs, a variant of this, Restricted Boltzmann Machine (RBM) is designed to learn the probability distribution over input data but in a layered manner [rbm].
@@
In UL, competitive learning approaches, such as Self-Organizing Maps (SOMs) [som], each neuron competes to represent an input subset. Here a single neuron from a group of output neurons is activated while the other neurons adjust their individual values in regard to input data distribution. Generative Adversarial Networks (GAN) [gan] consist of any two networks, with one generating data (generative network) and the other judging the generated data (discriminating network). The prediction accuracy of the discriminating network is then used to evaluate the error for the generating network. This creates a form of competition between the discriminator and the generator to get better in their corresponding tasks. We can also use ensemble learning methods [ai-ens] comprising multiple learning methods for better performances.
@@
Generative Diffusion Models (GDMs), as introduced in [diff], represents a recent breakthrough in UL leveraging DL techniques, drawing inspiration from the principles of thermodynamic diffusion. GDMs have gained widespread recognition due to their remarkable ability to generate high-quality data and simple implementation procedures. In contrast to GANs, GDMs employ a denoising network that iteratively converges to an estimate of the real sample. This model works in two distinct phases: the forward and reverse diffusion processes [diff2]. In the forward diffusion phase, Gaussian noise is gradually introduced through a series of steps to create the target input for the denoising network. Subsequently, the denoising network is trained to reverse the noise effect for generating the original content.
@@ III-D3 Reinforcement Learning (RL)
In RL, an agent learns to behave in a particular environment by performing thousands of actions and getting rewards or penalties based on those actions [ai-rl]. This behavior (formally known as policy) is defined by the set of actions the agent learns from its experiences. The environment is defined by some mathematical models, the most common one is the Markov Decision Process (MDP) [markov]. Here the feedback is neither provided using explicit labels like in SL nor the model is learned like in UL, but the behavior of an agent is learned through the rewards or penalties based on the set of actions taken by going from one state to another with a transition probability. The goal is to find out the optimum policy so that the total reward can be maximized (or the total penalty can be minimized) over a horizon of future time intervals given the current state of the agent. In Figure 11, we show a generic structure of an RL framework as an MDP model.
@@
ML Approaches: Depending on whether an RL model is explicitly created or not, RL can be fundamentally divided into two major categories: model-based and model-free RL methods [rl-intro]. In the model-based methods, the transition probabilities between different states are assumed to be known, whereas, in model-free methods, these probabilities are learned through iterations. Dynamic Programming (DP) [dp] is the most popular model-based method used in practice. However, model-free methods like the Monte Carlo (MC) method [mc] are most commonly used due to their flexibility and practicality in real systems. Q-learning [q] is one of the most popular model-free methods where ”Q” refers to the expected rewards for an action taken in a given state over the time horizon, known as the value function. Another important counterpart of Q-learning is the State-Action-Reward-State-Action (SARSA) learning method where the agent learns the optimal policy in an online fashion [sarsa]. This Q-learning is extended to the context of stochastic games [sg] involving multiple agents in [marl], which is also known as the Multi-Agent Reinforcement Learning (MARL) method. Another approach for model-free RL is to learn the policy directly instead of learning the value functions, which is known as the Policy Gradient (PG) method [pg]. If the policy gradient is estimated in a deterministic fashion, it is called the Deterministic Policy Gradient (DPG) method [dpg]. To have the benefits of both approaches, the Actor-Critic (AC) method is proposed in [ac] where the critic estimates the value function and the actor updates the policy gradient in the direction suggested by the critic.
@@
DL Approaches: General RL approaches without NNs can work well for small-scale experiments. However, when the action or state space is really large, the computation complexity exponentially increases. This phenomenon is quite common in practical systems like communication networks. To estimate the value functions or policies with RL frameworks with large state or action spaces, DL approaches can be very useful. This learning approach is also known as Deep Reinforcement Learning (DRL). The most popular and simplest DRL approach is the Deep Q-Network (DQN) [dqn] method. In this approach, instead of an iterative approach for updating Q values in the Q-table, an NN is used to estimate the Q function value approximately. To prevent a large overestimation of action values, another DL framework is introduced on top of DQN for a fair evaluation of policies in Double Deep-Q-Network (DDQN) [ddqn, ddqn2]. In Dueling Deep-Q-Network [dueldqn] method, both state and action values are separately estimated. As the expected value function may be overestimated as the expected value does not capture the complete probability distribution of random variables, Distributional Deep-Q-Network [disdqn] is considered to update the Q function value based on its distribution. In the case of continuous action spaces, DL aided DPG method, known as Deep Deterministic Policy Gradient (DDPG) Q-Learning [ddpg] provides better results. To deal with partial observable environments, Deep Recurrent Q-Networks (DRQN) [drqn] by introducing an LSTM layer in the FCNN architecture of DQN. Similarly, as for RL, MARL approaches can be efficiently solved using DRL architecture for each agent, namely known as a Multi-Agent Deep Reinforcement Learning (MADRL) framework.